A machine elastic bands with breaking tension normally distributed with mean 45 N and
standard deviation 4.36 N.
a. In order to test whether there is a change in the mean breaking tension, a random
sample of 50 was tested and found to have a mean breaking tension of 43.46 N.
i. Determine a 95% confidence interval for the population mean breaking tension
based on the sample mean assuming an unchanged standard deviation.
ii. Make a conclusion for answer (a)(i).
b. If the standard deviation has changed to sigma, find the least value of sigma
for a 95 % confidence interval for the population mean to contain 45 N.
1
Expert's answer
2020-04-22T18:33:18-0400

a. We need to construct the 95% confidence interval for the population "\mu." The following information is provided: "\bar{X}=43.46,\sigma=4.36,n=50,\alpha=0.05"

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96." The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-z_c\times{\sigma \over \sqrt{n}}, \bar{X}+z_c\times{\sigma \over \sqrt{n}})="

"=(43.46-1.96\times{4.36 \over \sqrt{50}}, 43.46+1.96\times{4.36 \over \sqrt{50}})\approx"

"\approx(42.251,44.669)"



b. Therefore, based on the data provided, the 95% confidence interval for the population mean is "42.251<\mu<44.669," which indicates that we are 95% confident that the true population mean "\mu"

 is contained by the interval "(42.251,44.669)."


c.


"\bar{X}+z_c\times{\sigma \over \sqrt{n}}\geq45"

"43.46+1.96\times{\sigma \over \sqrt{50}}\geq45"

"\sigma\geq{45-43.46 \over 1.96}\times\sqrt{50}"

"\sigma\geq5.555839"

"\sigma\approx5.56"